# Follow-Your-Pose v2: Multiple-Condition Guided Character Image Animation for Stable Pose Control

### v1

@@ 2.1. Diffusion Models for Video Generation
Compared with aforementioned GAN-based methods  (Mirza and Osindero, 2014; Siarohin et al., 2021; Zhao and Zhang, 2022), diffusion models (Song et al., 2020; Khachatryan et al., 2023; Guo et al., 2023a; Chen et al., 2023, 2024; Zhang et al., 2023b; Wu et al., 2023; Molad et al., 2023; Esser et al., 2023; Guo et al., 2023b; Ma et al., 2023; He et al., 2021; Ye et al., 2023; Zhang et al., 2023a) excel in generating high-quality contents and exhibit greater training efficiency. Diffusion models were first proposed to generate images (Ho et al., 2020; Song et al., 2020). Many studies adapt image synthesis pipelines for video generation (Khachatryan et al., 2023; Guo et al., 2023a; Chen et al., 2023, 2024; Wu et al., 2023). Text2Video-Zero (Khachatryan et al., 2023) incorporates motion dynamics in generated frames via cross-frame attention in zero-shot style. Animatediff (Guo et al., 2023a) finetunes a plug-and-play motion module that can be integrated into any text-to-image models to obtain animations in a personalized style. VideoCrafter (Chen et al., 2023, 2024) achieves both text-to-video and image-to-video generation and produces videos of high resolution. Besides text, I2V-Adapter (Guo et al., 2023b) extracts semantics from images to condition T2V generation.
@@ 2.2. Pose-Controllable Character Animation
With the development of diffusion model (Ho et al., 2020), existing methods incorporate pose sequences to guide the generation of the character video clips (Zhang et al., 2023b; Ma et al., 2024; Karras et al., 2023; Wang et al., 2023; Chang et al., 2023; Xu et al., 2023, 2023; Hu et al., 2023). DreamPose (Karras et al., 2023) proposes a finetuning framework to adapt the Stable Diffusion model (Rombach et al., 2022) to a pose-and-image guided video synthesis model. AnimateAnyone (Hu et al., 2023) designs ReferenceNet to merge detailed human body features, ensuring character appearance consistency, while employing a pose guider and temporal modeling for controllable and continuous movements. To enhance the faithfulness, generalizability, and compositionality of character animation, DisCo (Wang et al., 2023) disentangles the control conditions (i.e., character foreground, background, and poses) by introducing multiple ControlNets for different feature embeddings and proposes a human attribute pre-training framework. Moreover, a two-stage training scheme (Ma et al., 2024) is proposed to address the lack of comprehensive paired video-pose datasets, leveraging available image pose pairs, pose-free videos, and a pre-trained text-to-image model to generate pose-controllable character videos. However, both GAN-based and diffusion-based methods fail to effectively address the issue of background stability and occlusion among body parts.

### v3

@@ 2.1. Diffusion Models for Video Generation
Compared with aforementioned GAN-based methods  (Mirza and Osindero, 2014; Siarohin et al., 2021; Zhao and Zhang, 2022), diffusion models (Song et al., 2020; Khachatryan et al., 2023; Guo et al., 2023a; Chen et al., 2023, 2024; Zhang et al., 2023b; Wu et al., 2023; Molad et al., 2023; Esser et al., 2023; Guo et al., 2023b; Ma et al., 2023; He et al., 2021; Ye et al., 2023; Zhang et al., 2023a; Ma et al., 2024c, b) excel in generating high-quality contents and exhibit greater training efficiency. Diffusion models were first proposed to generate images (Ho et al., 2020; Song et al., 2020). Many studies adapt image synthesis pipelines for video generation (Khachatryan et al., 2023; Guo et al., 2023a; Chen et al., 2023, 2024; Wu et al., 2023). Text2Video-Zero (Khachatryan et al., 2023) incorporates motion dynamics in generated frames via cross-frame attention in zero-shot style. Animatediff (Guo et al., 2023a) finetunes a plug-and-play motion module that can be integrated into any text-to-image models to obtain animations in a personalized style. VideoCrafter (Chen et al., 2023, 2024) achieves both text-to-video and image-to-video generation and produces videos of high resolution. Besides text, I2V-Adapter (Guo et al., 2023b) extracts semantics from images to condition T2V generation.
@@ 2.2. Pose-Controllable Character Animation
With the development of diffusion model (Ho et al., 2020), existing methods incorporate pose sequences to guide the generation of the character video clips (Zhang et al., 2023b; Ma et al., 2024a; Karras et al., 2023; Wang et al., 2023; Chang et al., 2023; Xu et al., 2023, 2023; Hu et al., 2023). DreamPose (Karras et al., 2023) proposes a finetuning framework to adapt the Stable Diffusion model (Rombach et al., 2022) to a pose-and-image guided video synthesis model. AnimateAnyone (Hu et al., 2023) designs ReferenceNet to merge detailed human body features, ensuring character appearance consistency, while employing a pose guider and temporal modeling for controllable and continuous movements. To enhance the faithfulness, generalizability, and compositionality of character animation, DisCo (Wang et al., 2023) disentangles the control conditions (i.e., character foreground, background, and poses) by introducing multiple ControlNets for different feature embeddings and proposes a human attribute pre-training framework. Moreover, a two-stage training scheme (Ma et al., 2024a) is proposed to address the lack of comprehensive paired video-pose datasets, leveraging available image pose pairs, pose-free videos, and a pre-trained text-to-image model to generate pose-controllable character videos. However, both GAN-based and diffusion-based methods fail to effectively address the issue of background stability and occlusion among body parts.